Bob paid $3 less than Tim at a local pizza parlor. Together they spent $7.80. How much did Bob spend at the local pizza parlor?
Answer by nerdybill(7384)

Bob paid $3 less than Tim at a local pizza parlor. Together they spent $7.80. How much did Bob spend at the local pizza parlor?
.
Let T = amount Tim paid
T-3 = amount Bob paid
.
From: "Together they spent $7.80" we get our equation:
T + (T-3) = 7.80
2T - 3 = 7.80
2T = 10.80
T = 5.40 (amount Tim paid)
.
amount Bob paid:
T-3 = 5.40-3 = $2.40